OpenStudy (anonymous):

1. A football team plays in a city that has a fair chance of rain on any given day. Suppose this upcoming Sunday there is a 62% chance of rain. If it is raining, they can expect to score 3 touchdowns. If the weather is clear, they can expect to score 5 touchdowns. What is the expected number of touchdowns they score? 2. There is a prize drawing for a new computer and a new stereo system.Suppose there are 1500 raffle tickets sold. The computer is worth $1000 and the stereo system is worth $500. If the tickets cost $5, what is the expected value if you purchase one ticket?

OpenStudy (anonymous):

\[.62\times 3+.48\times 5\]

OpenStudy (anonymous):

I had 3.76

OpenStudy (anonymous):

Isn't it .38?

OpenStudy (anonymous):

second one we can work as follows, suppose you by all the raffle tickets you spend \[1500\times \$5=\$7500\] you get back $1500 for a net loss of $6000 divided by the total number of tickets the expected loss per ticket is \[\$6000\div 1500=\$4\] so expected value is -$4

OpenStudy (anonymous):

Satellite, could you please help me once you're done here? :)

OpenStudy (anonymous):

i get for the first one \[.62\times 3+.38\times 5=3.76\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yw